References

[1] A. V. Anisovich, V. Burkert, M. Dugger, E. Klempt, V. A. Nikonov, B. G. Ritchie, A. Sarantsev and U. Thoma, Proton-η′ interactions at threshold (2018)

[2] M. C. Kunkel et al., Exclusive photoproduction of π^{0} up to large values of Mandelstam variables s, t and u with CLAS, Phys. Rev. C98 (2018) (1), p. 015207

[3] P. Roy et al., Measurement of the beam asymmetry Σ and the target asymmetry T in the photoproduction of ω mesons off the proton using CLAS at Jefferson Laboratory, Phys. Rev. C97 (2018) (5), p. 055202

[4] Z. Akbar et al., Measurement of the helicity asymmetry E in ω → π^{+}π^{−}π^{0} photoproduction, Phys. Rev. C96 (2017) (6), p. 065209

[5] S. Adhikari et al., Strange Hadron Spectroscopy with a Secondary KL Beam at GlueX (2017)

[6] A. V. Anisovich, V. Burkert, P. M. Collins, M. Dugger, E. Klempt, V. A. Nikonov, B. G. Ritchie, A. V. Sarantsev and U. Thoma, N^{∗} → Nη′ decays from photoproduction of η′-mesons off protons, Phys. Lett. B772 (2017), pp. 247–252

[7] P. Collins et al., Photon beam asymmetry Σ in the reaction γ⃗p → pω for E_{γ} = 1.152 to 1.876 GeV, Phys. Lett. B773 (2017), pp. 112–120

[8] P. T. Mattione et al., Differential cross section measurements for γn → π^{−}p above the first nucleon resonance region, Phys. Rev. C96 (2017) (3), p. 035204

[9] A. V. Anisovich et al., Differential cross sections and polarization observables from CLAS K* photoproduction and the search for new N* states, Phys. Lett. B771 (2017), pp. 142–150

[10] M. Dugger et al., Design and construction of a high-energy photon polarimeter, Nucl. Instrum. Meth. A867 (2017), pp. 115–127

[11] P. Collins et al., Photon beam asymmetry Σ for η and η′ photoproduction from the proton, Phys. Lett. B771 (2017), pp. 213–221

[12] H. Al Ghoul et al., Measurement of the beam asymmetry Σ for π^{0} and η photoproduction on the proton at E_{γ} = 9 GeV, Phys. Rev. C95 (2017) (4), p. 042201

[13] C. A. Paterson et al., Photoproduction of Λ and Σ^{0} hyperons using linearly polarized photons, Phys. Rev. C93 (2016) (6), p. 065201

[14] H. Al Ghoul et al., First results from the GlueX experiment, AIP Conf. Proc. 1735 (2016), p. 020001

\underline{Institution}  \underline{Major/Conc.}  \underline{Degree, Year}
Undergraduate: Northern Arizona State University Physics/Mathematics B.S., 1993
Graduate:  Arizona State University  Physics  Ph.D., 2001
Postdoctoral: Arizona State University  Physics  2002-2006
Associate Professor (tenure-track) Arizona State University 2017 to present
Associate Research Professor  Arizona State University 2013 to 2017
Assistant Research Professor  Arizona State University 2006-2013
Postdoctoral Research Associate Arizona State University 2002-2006

0.1 Michael R. Dugger

Professional preparation Appointments Invited talks over the last three years

“Overview of Spectroscopy Results in Meson Photoproduction with Polarization Observables,” HADRON, Newport News, VA., Septrmber 14, 2015.

Publications over the last three years

1. A. V. Anisovich, V. Burkert, M. Dugger, E. Klempt, V. A. Nikonov, B. G. Ritchie, A. Sarantsev and U. Thoma, Proton-η′ interactions at threshold (2018)

2. M. C. Kunkel et al., Exclusive photoproduction of π^{0} up to large values of Mandelstam variables s, t and u with CLAS, Phys. Rev. C98 (2018) (1), p. 015207

3. P. Roy et al., Measurement of the beam asymmetry Σ and the target asymmetry T in the photoproduction of ω mesons off the proton using CLAS at Jefferson Laboratory, Phys. Rev. C97 (2018) (5), p. 055202

4. Z. Akbar et al., Measurement of the helicity asymmetry E in ω → π^{+}π^{−}π^{0} photoproduction, Phys. Rev. C96 (2017) (6), p. 065209

5. S. Adhikari et al., Strange Hadron Spectroscopy with a Secondary KL Beam at GlueX (2017)

6. A. V. Anisovich, V. Burkert, P. M. Collins, M. Dugger, E. Klempt, V. A. Nikonov, B. G. Ritchie, A. V. Sarantsev and U. Thoma, N^{∗} → Nη′ decays from photoproduction of η′-mesons off protons, Phys. Lett. B772 (2017), pp. 247–252

7. P. Collins et al., Photon beam asymmetry Σ in the reaction γ⃗p → pω for E_{γ} = 1.152 to 1.876 GeV, Phys. Lett. B773 (2017), pp. 112–120

8. P. T. Mattione et al., Differential cross section measurements for γn → π^{−}p above the first nucleon resonance region, Phys. Rev. C96 (2017) (3), p. 035204 9. A. V. Anisovich et al., Differential cross sections and polarization observables from CLAS

K* photoproduction and the search for new N* states, Phys. Lett. B771 (2017), pp. 142–150

10. M. Dugger et al., Design and construction of a high-energy photon polarimeter, Nucl. Instrum. Meth. A867 (2017), pp. 115–127

11. P. Collins et al., Photon beam asymmetry Σ for η and η′ photoproduction from the proton, Phys. Lett. B771 (2017), pp. 213–221

12. H. Al Ghoul et al., Measurement of the beam asymmetry Σ for π^{0} and η photoproduction on the proton at E_{γ} = 9 GeV, Phys. Rev. C95 (2017) (4), p. 042201

13. C. A. Paterson et al., Photoproduction of Λ and Σ^{0} hyperons using linearly polarized photons, Phys. Rev. C93 (2016) (6), p. 065201

14. H. Al Ghoul et al., First results from the GlueX experiment, AIP Conf. Proc. 1735 (2016), p. 020001